# Self-supervised 3D Point Cloud Completion via Multi-view Adversarial Learning

### v1

@@ IV-C2 Results on real-world scans
IV-C3 Comparison of computational complexity and model size
Table III provides a comprehensive comparison of the computational complexity, measured in Floating Point Operations (FLOPs), and the model size for our proposed method against several state-of-the-art approaches: USSPA, ACL-SPC, and P2C. Our model demonstrates a well-balanced approach in terms of both computational complexity and model size. With 15.10G FLOPs, it is considerably more efficient than USSPA. In terms of model size, our method, with 94.79M parameters, remains manageable and significantly smaller than both ACL-SPC and P2C. This balance makes our model an attractive option for practical applications where both computational resources and memory capacity are constrained.
@@ IV-C3 Comparison of computational complexity and model size
IV-D Ablation Study
IV-D1 Effects of on each module
To analyze the effect of each innovative component in our framework, including position encoding, curvature encoding, multi-view adversarial network, and density-aware radius estimation strategy, we conducted ablation experiments on synthetic data. As Table IV shows, the inclusion of all modules (PE, CE, MAN, and DARE) yields the best average performance, achieving the lowest average error of 9.08. In particular, adding only PE results in a noticeable reduction in the average error to 10.76, especially a substantial gain in the lamp category (from 24.22 to 16.89). We also compared the visualization results with and without using PE to visually observe the impact of PE on the reconstruction process. As Figure 6 shows, without PE, the model predicts some inaccurate points that significantly deviate from the partial input and the ground truth. However, after incorporating PE, these inaccurate points are noticeably reduced. We can conclude from both the quantitative and qualitative results that PE is able to effectively utilize the positional relationships among partial points to infer more accurate point positions.
Additionally, it is evident from Table IV that incorporating both PE and CE leads to a further reduction in the average error to 9.87. CE appears to complement PE by providing additional curvature information. The combination significantly improves performance in categories like table, couch, and cabinet. Moreover, as observed from the visual results in Figure 7, after embedding CE, when the surface of the incomplete input point cloud is compact and smooth, the corresponding reconstructed surface is also significantly smoother. These demonstrate that CE can effectively learn the curvature features of the partial input and leverage them to reconstruct more realistic surfaces.
In addition to PE and CE, the inclusion of MAN further enhances the performance as demonstrated in Table IV, with an average improvement of 0.55. Note that the density-aware radius estimation algorithm also contributes positively to the model’s performance, with significant improvements in categories like chair and table.
@@ IV-D2 Ablation on supervisory signals
To evaluate the impact of these two supervisory signals on the model’s performance, we conducted ablation studies on the synthetic dataset. We used our complete model and performed experiments by separately removing $L_{part}$ and $L_{rend}$. The results show that non-converge occurs in the absence of either $L_{part}$ or $L_{rend}$. For the case of lacking $L_{rend}$, the CD values achieved on the cabinet, car, couch, and table categories are 36.56, 18.19, 19.68, and 74.83, respectively, while convergence is not achieved on other categories. On the other hand, when $L_{part}$ is removed, CD values on cabinet, chair, car, couch, table, and watercraft categories are 29.13, 21.12, 7.36, 12.00, 12.53, and 9.78, respectively, while convergence failed for plane and lamp categories. We believe that the reason for the non-convergence is twofold. First, the training of GAN itself is inherently unstable [62, 69]. Second, the inclusion of DARE makes training more challenging. This is because DARE requires a relatively accurate initial depth map rendered from the predicted results before estimating a more precise point radius. If the initial depth map has significant deviations, the estimated point radius will also deviate accordingly, leading to significant errors in the second-round rendering. Consequently, the GAN module might collapse, ultimately preventing the model from converging. The lack of either $L_{rend}$ or $L_{part}$, which provides direct supervision, increases the difficulty for the model to predict accurate shapes, contributing to the non-convergence. However, based on the results from the converged categories, $L_{rend}$ exhibits stronger supervisory capabilities compared to $L_{part}$.
V Conclusion
@@
In the scenario of self-supervised point cloud completion, it is critical to fully utilize the partial observations in the training data as priors due to the lack of complete shapes as supervision. To this end, we have presented a novel approach that effectively exploits the objects’ intrinsic object-level and category-specific geometric similarity to infer complete shapes without the need for multi-view partial scans and masked operation. Our method demonstrates strong capabilities in both inferring complete shapes and preserving incomplete input, resulting in reconstructed shapes that are both reasonable and accurate, with minimal noise. Quantitatively and visually, our approach significantly surpasses current state-of-the-art self-supervised methods, as well as a portion of unpaired methods, by a large margin. In future work, we aim to extend our work to large-scale indoor and outdoor scene completion tasks. We hope this work will inspire future research in 3D object and scene completion, reconstruction, and understanding tasks.

### v2

@@ IV-D1 Effects of each module
To analyze the effect of each innovative component in our framework, including position encoding, curvature encoding, multi-view adversarial network, and density-aware radius estimation strategy, we conduct ablation experiments on synthetic data. As Table IX shows, the inclusion of all modules (PE, CE, MAN, and DARE) yields the best average performance, achieving the lowest average error of 9.08. In particular, adding only PE results in a noticeable reduction in the average error to 10.76, especially a substantial gain in the lamp category (from 24.22 to 16.89). We also compared the visualization results with and without using PE to observe the impact of PE on the reconstruction process. As Figure 8 illustrates, without PE, the model predicts some inaccurate points, including outliers that significantly deviate from the partial input and the ground truth. However, after incorporating PE, these inaccurate points are noticeably reduced. We attribute this effect to the following reasons: PE captures the local patterns of the coarse shape $P_{c}$ and the partial input $P_{in}$. As discussed in Section III-A, the PRN retrieves the most relevant PE from $P_{in}$ for $P_{c}$ and uses them to predict an offset for $P_{c}$. Consequently, even though the local regions in $P_{c}$ may be noisy, the PRN leverages the structured and noise-free local patterns from $P_{in}$ to calibrate the positions of noisy local points in $P_{c}$.
@@
Additionally, it is evident from Table IX that incorporating both PE and CE reduces the average CD further to 9.87. CE complements PE by providing additional curvature information, significantly improving performance in categories such as table, couch, and cabinet. Moreover, as observed from the visual results in Figure 9, embedding CE results in a noticeably smoother reconstructed surface when the incomplete input point cloud surface is compact and smooth. Like PE, CE encodes the curvature patterns of the local surfaces of $P_{c}$ and $P_{in}$. Through attention-based retrieval, PRN leverages the most relevant smooth surface knowledge from $P_{in}$ to refine the rough local surfaces of $P_{c}$.
In addition to PE and CE, the inclusion of MAN further enhances the performance as demonstrated in Table IX, with an average improvement of 0.55. Note that the density-aware radius estimation algorithm also contributes positively to the model’s performance, with significant improvements in categories like chair and table.
@@ IV-D2 Ablation on supervisory signals
To evaluate the impact of these two supervisory signals on the model’s performance, we conduct ablation studies on the synthetic dataset. We use our complete model and perform experiments by separately removing $L_{part}$ and $L_{rend}$. The results show that non-converge occurs in the absence of either $L_{part}$ or $L_{rend}$. For the case of lacking $L_{rend}$, the CD values achieved on the cabinet, car, couch, and table categories are 36.56, 18.19, 19.68, and 74.83, respectively, while convergence is not achieved on other categories. On the other hand, when $L_{part}$ is removed, CD values on cabinet, chair, car, couch, table, and watercraft categories are 29.13, 21.12, 7.36, 12.00, 12.53, and 9.78, respectively, while convergence failed for plane and lamp categories. We believe that the reason for the non-convergence is twofold. First, the training of GAN itself is inherently unstable [64, 71]. Second, the inclusion of DARE makes training more challenging. This is because DARE requires a relatively accurate initial depth map rendered from the predicted results before estimating a more precise point radius. If the initial depth map has significant deviations, the estimated point radius will also deviate accordingly, leading to significant errors in the second-round rendering. Consequently, the GAN module might collapse, ultimately preventing the model from converging. The lack of either $L_{rend}$ or $L_{part}$, which provides direct supervision, increases the difficulty for the model to predict accurate shapes, contributing to the non-convergence. However, based on the results from the converged categories, $L_{rend}$ exhibits stronger supervisory capabilities compared to $L_{part}$.
V Limitation and Future Work
Although our method achieves outstanding results, it also has some limitations. First, since MAL-SPC relies on the geometric similarity of objects within the same category, it requires category information of the partial observations. Besides, our model performs poorly on unseen categories and rare objects, despite outperforming other self-supervised approaches. In the future, we will attempt to eliminate the dependency on category information using methods such as clustering and contrastive learning. Additionally, we plan to leverage large quantities of 2D images and current Large Vision Models to provide 2D priors for partial point clouds, so as to improve the generalization capability of self-supervised methods.
VI Conclusion
@@
In the scenario of self-supervised point cloud completion, it is critical to fully utilize the prior geometries from partial observations due to the lack of complete shapes as supervision. To this end, we have presented a novel approach that effectively exploits the objects’ intrinsic region-level and category-specific geometric similarity to infer complete shapes without the need for multi-view partial scans and masked operation. Our method demonstrates strong capabilities in both inferring complete shapes and preserving incomplete input, resulting in reconstructed shapes that are both reasonable and accurate, with minimal noise. Quantitatively and visually, our approach surpasses current state-of-the-art self-supervised methods, as well as a portion of unpaired methods. We hope this work inspires future research in 3D object and scene completion, reconstruction, and understanding tasks.